# MergeMix: A Unified Augmentation Paradigm for Visual and Multi-Modal Understanding

### v2

@@ 1 Introduction
Multi-modal Large Language Models (MLLMs) (liu2024llava; bai2025qwen2.5; tong2024cambrian) have recently demonstrated remarkable capabilities in integrating visual and textual information, enabling a wide range of applications from visual question answering to multi-modal reasoning. Since MLLMs are typically pre-trained on massive web-scale datasets, forcing them to possess a wide range of knowledge and general reasoning capabilities, Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL)-based preference optimization (yang2025visionthink) have emerged as two primary paradigms for aligning MLLMs with human preferences and specific task requirements. However, SFT depends on high-quality instruction–response annotations and optimizes the likelihood of reference responses, which does not explicitly model relative preferences between outputs. RL-based methods such as RLHF are more preference-aware, but require an additional reward model that may introduce bias or be exploited by the reward signal.
Due to the shortcomings of data quality and training efficiency, some works (zhu2024seva; zhu2025nsft; luo2024probing; tan2025beyond; wang2024sima) try to build performance pairs for optimization. How to build the preference pair with control and high-quality data for model training is the remaining open question. For example, SeVa (zhu2024seva) proposed a preference optimization method by building a loser through some classic augmentation (i.e., RandomCrop). Then, select the different responses for optimizing the model by a DPO loss (rafailov2023dpo). However, these methods have two drawbacks: the augmentations are highly random, and the DPO loss cannot be related to the data, which means SeVa can only select useful training data. Those technical causes SeVa can not control the quality of the loser, which is harmful for some visual question answering tasks, and reduces the training data by selecting “hard negatives”. Hence, we investigate an interesting question: Is it necessary to propose novel techniques rather than some classical machine learning methods in the MLLM scenario?
@@
Figure 1: Efficiency and for MergeMix: (a) The training time vs. accuracy of mixup methods with the DeiT-Small model. (b) The image classification Top-1 accuracy vs. training epochs of different mixup methods on the CIFAR100 dataset with the DeiT-Tiny model. (c) The radar plot of the results on part VQA tasks by LLaVA-7B, LLaVA with SFT, and MergeMix.
@@ 2 Related Work
The Mixup method mitigates model overfitting by generating augmented samples through mixing two different images within a mini-batch. Broadly, Mixup methods can be categorized into two types: Static, which relies on human priors or randomness, and Adaptive, a data-dependent type that leverages certain metrics to guide the mixing process. (i) Static: MixUp (zhang2017mixup) generates mixed samples via linear interpolation with $\lambda$. CutMix (yun2019cutmix) extends this idea from the global pixel level to a local patch level by constructing a mask of size proportional to $\lambda$ to mix images. ResizeMix (qin2020resizemix) ensures that features from at least one class are always preserved in the mixed sample by resizing the source image before mixing. Other methods, e.g., FMix (harris2020fmix), SmoothMix (lee2020smoothmix), GridMix (Baek2021gridmix), and StarMix (jin2024starlknet), focus on improving the mask to obtain more suitable mixed samples. (ii) Adaptive: SaliencyMix (uddinsaliencymix) employs a saliency extractor to identify informative patches in images for mixing. Attentive-CutMix (icassp2020attentive) and SuperMix (dabouei2021supermix) utilize a teacher model to guide mask generation. PuzzleMix (kim2020puzzle) and Co-Mix (kim2021co) generate appropriate masks based on gradient information obtained from a forward pass of the samples. AutoMix (liu2022automix) and AdAutoMix (qin2024adversarial) adopt an end-to-end bi-optimization paradigm to produce mixed samples. TransMix (chen2022transmix), SMMix (chen2023smmix), and MixPro (zhao2023mixpro) specifically enhance ViTs by computing attention scores from a forward pass to generate feature-aware masks and further refine the label ratio through attention scores. DiffuseMix (islam2024diffusemix) and GenMix (islam2024genmix) generate mixed samples by a diffusion model (zhu2025obs) for label preserving.
@@
Token compression methods (bolya2023tome; tao2025dycoke; shao2025holitom; chen2025streamingtom; tao2025omnizip) mainly propose to merge or drop redundant tokens to achieve efficiency and acceleration. In MLLMs, images and texts will incur a significant number of tokens, which are often full of redundant information. Obvious researchers bring the token compression into MLLMs. Overall, we divide the methods that reduce tokens into 2 types, Reduce in Encoder and Reduce in Decoder. (i) Reduce in Encoder: MADTP (cao2024madtp) aims to achieve MLLM acceleration by purging visual tokens. LLaVA-PruMerge (shang2024llavapruerge) uses the attention of [CLS] token to select clustering centers and then merges the remaining tokens with lower attention through a KNN clustering and weighted clustering center updating mechanism. VisionZip (yang2025visionzip), instead, retains visual tokens with high attention scores and subsequently merges the remaining tokens through clustering. Others, such as TokenPacker (li2025tokenpacker), AVG-LLaVA (lan2025avg), MustDrop (liu2024mustdrop), and LLaVolta (chen2024LLaVolta), achieve acceleration by choosing a metric to sample TopK visual tokens. FastVLM (vasu2025fastvlm) proposes an Efficient Vision Encoder to reduce visual tokens. (ii) Reduce in Deocder: PyramidDrop (xing2024pyramiddrop) divides the token compression process in LLM into multiple stages, which employs a pyramidal token drop to avoid losing too much visual information in shallower layers. ATP-LLaVA (ye2025atp) proposes an Adaptive Token Pruning (ATP) module that reduces the number of tokens in the decoder layer. ZipVL (he2024zipvl) proposes a dynamic ratio allocation strategy via the importance token, adaptively determined based on the distribution of attention scores in a particular layer, rather than a fixed hyperparameter.
@@ 3 Preliminaries
where $(x,y^{+})$ and $(x,y^{-})$ denote the preferred and dispreferred responses respectively, and $s_{\theta}$ is a scoring function that reflects model preference. Different approaches (e.g., PPO (schulman2017ppo), DPO (rafailov2023dpo)) instantiate $\pi_{\theta}$ in various ways, but all share the same principle.
@@
where $\pi_{\theta}$ denotes the policy model (in our case, an MLLM such as LLaVA-v1.5 (liu2024llava) or Qwen-VL (bai2025qwen2.5)), and $\pi_{\text{ref}}$ represents a frozen reference model used to preserve alignment with the original pre-trained distribution. $\sigma$ is the sigmoid function, and $\beta>0$ is a temperature-like scaling factor that controls the sharpness of preference separation. Intuitively, DPO encourages the policy to assign a higher likelihood to preferred responses ($y^{+}$) than to non-preferred ones ($y^{-}$), while maintaining proximity to the reference model.
Figure 2: The overall of the two scenarios of MergeMix: (a) MergeMix for Image Classification: The image is processed by the ToMe encoder, with Attention Score Recovery and TopK sampling to generate the corresponding class prediction. (b) MergeMix for MLLM: Preference pairs are encoded by the vision model with token merging, and the LLM decoder generates response text for the loser and winner, optimized via a ranking loss.
@@ 4.1 Image Mixing via Token Merge
Since token merging aggregates non-similar tokens into compact representations $Z_{K}$, the resulting matrix preserves local feature structures more effectively. In contrast, the vanilla TopK selection adopts a greedy sampling strategy with linear complexity $\mathcal{O}(N)$, which discards low-ranked tokens directly and thus loses spatial relationships. Alternatively, the Bipartite Soft Matching (BSM) approach performs global pairwise matching with quadratic complexity, yielding a more balanced and globally optimal merging of tokens. To reconstruct the full-resolution attention map, we introduce a recovering function $\mathcal{R}_{K\to L}(\cdot,\cdot)$ that expands the merged attention map $A_{K}$ back to its original length $A_{L}$, which shows in Fig 5 and according to Eq. (6):
@@ 4.2 A Unified Augmentation Paradigm: From Image Classification to MLLMs
Under this optimization objective, the role of the mixing ratio $\lambda$ is to serve as a metric that quantifies the presence of feature information from the two samples. While this metric cannot directly reflect the true characteristics of the data, certain adaptive methods can constrain the model to generate mixed samples where the mixing ratio progressively approximates the target value (jin2024survey). Some works, like LUMix (sun2022lumix), DecoupleMix (liu2023decoupledmix), and SUMix (qin2024sumix), use a defined policy for some hand-crafted mixup methods, and find that it is more efficient than optimizing a better mask way.
@@
LLaVA (liu2024llava) uses a standard conditional language modeling loss for SFT. In MLLMs, we are given an instruction–response pair $(x,y)$, where $x$ denotes the multi-modal data, inducing vision and text, and $y=(y_{1},y_{|y|})$ denotes the target response. The SFT loss is defined as:
@@
This objective needs to maximize the likelihood of GT responses, aiming to align the data. In Section 3, we introduced the DPO loss, which can be decomposed into two components: the SFT part and the ranking optimization part. In our approach, we replace the ranking component with SimPO (meng2024simpo), where $y$ denotes the target sequence (response) and $|y|$ denotes its length. Furthermore, since $\lambda$ reflects information similarity between augmented and raw image (interpreted as “loser degree” in MLLMs), we link it to $\gamma\to 1-\hat{\lambda}$: larger $\lambda$ represent higher similarity and harder discrimination, reduces $\gamma$ to avoid over-optimization on trivial differences; smaller $\lambda$ represent greater dissimilarity and easier tasks increases $\gamma$ to strengthen constraints for clearer preference distinction. The mixed SimPO loss replacement is Eq. (11):
$)].\displaystyle\mathcal{L}^{\text{Mix}}_{\mathrm{SimPO}}=-\mathbb{E}_{(x,\hat{x},y)\sim\mathcal{D}}\Big[\log\sigma\big(\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)-\frac{\beta}{|y|}\log\pi_{\theta}(y\mid\hat{x})-(1-\hat{\lambda}$$\big)\Big$]. (11)
@@ 5 Experiments
To evaluate the performance of MergeMix, we compared with some mainstream mixup methods, i.e. Mixup (zhang2017mixup), CutMix (yun2019cutmix), FMix (harris2020fmix), SmoothMix (lee2020smoothmix), GridMix (Baek2021gridmix), ResizeMix (qin2020resizemix), SaliencyMix (uddinsaliencymix), Attentive-CutMix (icassp2020attentive), PuzzleMix (kim2020puzzle), GuidedMixup (kang2023guidedmixup), AutoMix (liu2022automix) and AdAutoMix (qin2024adversarial). DeiT (icml2021deit), TransMix (chen2022transmix), SMMix (chen2023smmix), MixPro (zhao2023mixpro) and TdAttenMix (wang2025tdattenmix) for some ViT-based methods. The training configures about datasets and methods follows the open-source library OpenMixup (2022openmixup).
@@
To evaluate the training paradigm that we proposed, we compare with three different system-level baselines: (1) SFT with different training paradigms on LLaVA, including LLaVA-NeXT-7/13B (liu2024llavanext), SeVa-7B (zhu2024seva), SIMA (wang2024sima), and nSFT (zhu2025nsft). (2). Token reduction on LLaVA, including LLaVA-PruMerge+ (shang2024llavapruerge), VisionZip (yang2025visionzip), VisPrunner (zhang2024vispunner), VScan (zhang2025vscan), and LLaVA-Mini (zhang2025llava-mini). (3). RL training on Qwen2.5-VL-Instruction (bai2025qwen2.5), including VisionThink (yang2025visionthink).
For all classification results, we report the top-1 test accuracy in the last 10 training epochs for each trial. To facilitate comparison, we mark the best and second best results in bold and cyan. For the LLaVA benchmark, we use the LLaVA official code, and for the Qwen2.5-VL-Instruction benchmark. We use lmms-eval (zhang2024lmmsevalrealitycheckevaluation) for evaluation.
@@ 5.2 Datasets
In our paper, we mainly divided into 2 scenarios: Image Classification and MLLM Benchmark. The detailed information about datasets is described in Appendix B.1. For the image classification datasets, we choose 5 public classification datasets, including the small-scale dataset of CIFAR100 (krizhevsky2009learning), the large-scale dataset of ImageNet-1K (russakovsky2015imagenet), and the fine-grained datasets of CUB200 dataset (wah2011caltech), FGVC-Aircrafts dataset (maji2013fine), and Stanford-Cars dataset (jonathan2013cars). For the MLLM datasets, we choose 16 datasets, including visual question answering (VQAv2 (goyal2017vqav2), GQA (hudson2019gqa), VizWiz (gurari2018vizwiz), ScienceVQAI (lu2022scivqa), TextVQA (singh2019textvqa), MME-RealWorldQA (zhang2025realworldqa)), understanding (MME (Perception) (yin2023mme), MMBench (liu2025mmbench), $^{\text{CN}}, MMBench$$^{\text{CC}}$, POPE (F1 score) (li2023pope), SEEDI (li2023seed), MMStar (chen2024mmstar)), and reasoning (MMMU (yue2024mmmu), MMMU-Pro Standard (MMMU-Pros) (yue2024mmmu-pro), MathVista (lu2024mathvista)).

### v3

@@ 1 Introduction
Multi-modal Large Language Models (MLLMs) (Liu et al., 2024b; Bai et al., 2025; Tong et al., 2024) have recently demonstrated remarkable capabilities in integrating visual and textual information, enabling a wide range of applications from visual question answering to multi-modal reasoning. Since MLLMs are typically pre-trained on massive web-scale datasets, forcing them to possess a wide range of knowledge and general reasoning capabilities, Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL)-based preference optimization (Yang et al., 2025c) have emerged as two primary paradigms for aligning MLLMs with human preferences and specific task requirements. However, SFT depends on high-quality instruction–response annotations and optimizes the likelihood of reference responses, which does not explicitly model relative preferences between outputs. RL-based methods such as RLHF are more preference-aware, but require an additional reward model that may introduce bias or be exploited by the reward signal.
Due to the shortcomings of data quality and training efficiency, some works (Zhu et al., 2024; 2025b; Luo et al., 2024; Tan et al., 2025; Wang et al., 2024b) try to build performance pairs for optimization. How to build the preference pair with control and high-quality data for model training is the remaining open question. For example, SeVa (Zhu et al., 2024) proposed a preference optimization method by building a loser through some classic augmentation (i.e., RandomCrop). Then, select the different responses for optimizing the model by a DPO loss (Rafailov et al., 2023). However, these methods have two drawbacks: the augmentations are highly random, and the DPO loss cannot be related to the data, which means SeVa can only select useful training data. Those technical causes SeVa can not control the quality of the loser, which is harmful for some visual question answering tasks, and reduces the training data by selecting “hard negatives”. Hence, we investigate an interesting question: Is it necessary to propose novel techniques rather than some classical machine learning methods in the MLLM scenario?
@@
Figure 1: Efficiency and performance for MergeMix. (a) The training time vs. accuracy of mixup methods with the DeiT-Small model. (b) The image classification Top-1 accuracy vs. training epochs of different mixup methods on the CIFAR100 dataset with the DeiT-Tiny model. (c) The radar plot of the results on part VQA tasks by LLaVA-7B, LLaVA with SFT, and MergeMix.
@@ 2 Related Work
The Mixup method mitigates model overfitting by generating augmented samples through mixing two different images within a mini-batch. Broadly, Mixup methods can be categorized into two types: Static, which relies on human priors or randomness, and Adaptive, a data-dependent type that leverages certain metrics to guide the mixing process. (i) Static: MixUp (Zhang et al., 2017) generates mixed samples via linear interpolation with $\lambda$. CutMix (Yun et al., 2019) extends this idea from the global pixel level to a local patch level by constructing a mask of size proportional to $\lambda$ to mix images. ResizeMix (Qin et al., 2020) ensures that features from at least one class are always preserved in the mixed sample by resizing the source image before mixing. Other methods, e.g., FMix (Harris et al., 2020), SmoothMix (Lee et al., 2020), GridMix (Baek et al., 2021), and StarMix (Jin et al., 2024b), focus on improving the mask to obtain more suitable mixed samples. (ii) Adaptive: SaliencyMix (Uddin et al., 2021) employs a saliency extractor to identify informative patches in images for mixing. Attentive-CutMix (Walawalkar et al., 2020) and SuperMix (Dabouei et al., 2021) utilize a teacher model to guide mask generation. PuzzleMix (Kim et al., 2020) and Co-Mix (Kim et al., 2021) generate appropriate masks based on gradient information obtained from a forward pass of the samples. AutoMix (Liu et al., 2022) and AdAutoMix (Qin et al., 2024a) adopt an end-to-end bi-optimization paradigm to produce mixed samples. TransMix (Chen et al., 2022), SMMix (Chen et al., 2023), and MixPro (Zhao et al., 2023) specifically enhance ViTs by computing attention scores from a forward pass to generate feature-aware masks and further refine the label ratio through attention scores. DiffuseMix (Islam et al., 2024b) and GenMix (Islam et al., 2024a) generate mixed samples by a diffusion model (Zhu et al., 2025a) for label preserving.
@@
Token compression methods (Bolya et al., 2023; Tao et al., 2025a; Shao et al., 2025; Chen et al., 2025; Tao et al., 2025b) mainly propose to merge or drop redundant tokens to achieve efficiency and acceleration. In MLLMs, images and texts will incur a significant number of tokens, which are often full of redundant information. Obvious researchers bring the token compression into MLLMs. Overall, we divide the methods that reduce tokens into 2 types, Reduce in Encoder and Reduce in Decoder. (i) Reduce in Encoder: MADTP (Cao et al., 2024) aims to achieve MLLM acceleration by purging visual tokens. LLaVA-PruMerge (Shang et al., 2024) uses the attention of [CLS] token to select clustering centers and then merges the remaining tokens with lower attention through a KNN clustering and weighted clustering center updating mechanism. VisionZip (Yang et al., 2025b), instead, retains visual tokens with high attention scores and subsequently merges the remaining tokens through clustering. Others, such as TokenPacker (Li et al., 2025), AVG-LLaVA (Lan et al., 2025), MustDrop (Liu et al., 2024c), and LLaVolta (Chen et al., 2024a), achieve acceleration by choosing a metric to sample TopK visual tokens. FastVLM (Vasu et al., 2025) proposes an Efficient Vision Encoder to reduce visual tokens. (ii) Reduce in Deocder: PyramidDrop (Xing et al., 2024) divides the token compression process in LLM into multiple stages, which employs a pyramidal token drop to avoid losing too much visual information in shallower layers. ATP-LLaVA (Ye et al., 2025) proposes an Adaptive Token Pruning (ATP) module that reduces the number of tokens in the decoder layer. ZipVL (He et al., 2024) proposes a dynamic ratio allocation strategy via the importance token, adaptively determined based on the distribution of attention scores in a particular layer, rather than a fixed hyperparameter.
@@ 3 Preliminaries
where $(x,y^{+})$ and $(x,y^{-})$ denote the preferred and dispreferred responses respectively, and $s_{\theta}$ is a scoring function that reflects model preference. Different approaches (e.g., PPO (Schulman et al., 2017), DPO (Rafailov et al., 2023)) instantiate $\pi_{\theta}$ in various ways, but all share the same principle.
@@
where $\pi_{\theta}$ denotes the policy model (in our case, an MLLM such as LLaVA-v1.5 (Liu et al., 2024b) or Qwen-VL (Bai et al., 2025)), and $\pi_{\text{ref}}$ represents a frozen reference model used to preserve alignment with the original pre-trained distribution. $\sigma$ is the sigmoid function, and $\beta>0$ is a temperature-like scaling factor that controls the sharpness of preference separation. Intuitively, DPO encourages the policy to assign a higher likelihood to preferred responses ($y^{+}$) than to non-preferred ones ($y^{-}$), while maintaining proximity to the reference model.
Figure 2: Overview of MergeMix in two scenarios. (a) MergeMix for Image Classification: The image is processed by the ToMe encoder, with Attention Score Recovery and TopK sampling to generate the corresponding class prediction. (b) MergeMix for MLLM: Preference pairs are encoded by the vision model with token merging, and the LLM decoder generates response text for the loser and winner, optimized via a ranking loss.
@@ 4.1 Image Mixing via Token Merge
Since token merging aggregates non-similar tokens into compact representations $Z_{K}$, the resulting matrix preserves local feature structures more effectively. In contrast, the vanilla TopK selection adopts a greedy sampling strategy with linear complexity $\mathcal{O}(N)$, which discards low-ranked tokens directly and thus loses spatial relationships. Alternatively, the Bipartite Soft Matching (BSM) approach performs global pairwise matching with quadratic complexity, yielding a more balanced and globally optimal merging of tokens. To reconstruct the full-resolution attention map, we introduce a recovering function $\mathcal{R}_{K\to L}(\cdot,\cdot)$ that expands the merged attention map $A_{K}$ back to its original length $A_{L}$, which shows in Figure 5 and according to Eq. (6):
@@ 4.2 A Unified Augmentation Paradigm: From Image Classification to MLLMs
Under this optimization objective, the role of the mixing ratio $\lambda$ is to serve as a metric that quantifies the presence of feature information from the two samples. While this metric cannot directly reflect the true characteristics of the data, certain adaptive methods can constrain the model to generate mixed samples where the mixing ratio progressively approximates the target value (Jin et al., 2024a). Some works, like LUMix (Sun et al., 2022), DecoupleMix (Liu et al., 2023), and SUMix (Qin et al., 2024b), use a defined policy for some hand-crafted mixup methods, and find that it is more efficient than optimizing a better mask way.
@@
LLaVA (Liu et al., 2024b) uses a standard conditional language modeling loss for SFT. In MLLMs, we are given an instruction–response pair $(x,y)$, where $x$ denotes the multi-modal data, inducing vision and text, and $y=(y_{1},y_{|y|})$ denotes the target response. The SFT loss is defined as:
@@
This objective needs to maximize the likelihood of GT responses, aiming to align the data. In Section 3, we introduced the DPO loss, which can be decomposed into two components: the SFT part and the ranking optimization part. In our approach, we replace the ranking component with SimPO (Meng et al., 2024), where $y$ denotes the target sequence (response) and $|y|$ denotes its length. Furthermore, since $\lambda$ reflects information similarity between augmented and raw image (interpreted as “loser degree” in MLLMs), we link it to $\gamma\to 1-\hat{\lambda}$: larger $\lambda$ represent higher similarity and harder discrimination, reduces $\gamma$ to avoid over-optimization on trivial differences; smaller $\lambda$ represent greater dissimilarity and easier tasks increases $\gamma$ to strengthen constraints for clearer preference distinction. The mixed SimPO loss replacement is Eq. (11):
$)].\displaystyle\mathcal{L}^{\text{Mix}}_{\text{SimPO}}=-\mathbb{E}_{(x,\hat{x},y)\sim\mathcal{D}}\Big[\log\sigma\big(\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)-\frac{\beta}{|y|}\log\pi_{\theta}(y\mid\hat{x})-(1-\hat{\lambda}$$\big)\Big$]. (11)
@@ 5.1 State-of-the-art methods
To evaluate the performance of MergeMix, we compared with some mainstream mixup methods, i.e. Mixup (Zhang et al., 2017), CutMix (Yun et al., 2019), FMix (Harris et al., 2020), SmoothMix (Lee et al., 2020), GridMix (Baek et al., 2021), ResizeMix (Qin et al., 2020), SaliencyMix (Uddin et al., 2021), Attentive-CutMix (Walawalkar et al., 2020), PuzzleMix (Kim et al., 2020), GuidedMixup (Kang and Kim, 2023), AutoMix (Liu et al., 2022) and AdAutoMix (Qin et al., 2024a). DeiT (Touvron et al., 2021), TransMix (Chen et al., 2022), SMMix (Chen et al., 2023), MixPro (Zhao et al., 2023) and TdAttenMix (Wang et al., 2025) for some ViT-based methods. The training configures about datasets and methods follows the open-source library OpenMixup (Li et al., 2022).
@@
To evaluate the training paradigm that we proposed, we compare with three different system-level baselines: (1) SFT with different training paradigms on LLaVA, including LLaVA-NeXT-7/13B (Liu et al., 2024a), SeVa-7B (Zhu et al., 2024), SIMA (Wang et al., 2024b), and nSFT (Zhu et al., 2025b). (2). Token reduction on LLaVA, including LLaVA-PruMerge+ (Shang et al., 2024), VisionZip (Yang et al., 2025b), VisPrunner (Zhang et al., 2024b), VScan (Zhang et al., 2025a), and LLaVA-Mini (Zhang et al., 2025b). (3). RL training on Qwen2.5-VL-Instruction (Bai et al., 2025), including VisionThink (Yang et al., 2025c).
For all classification results, we report the top-1 test accuracy in the last 10 training epochs for each trial. To facilitate comparison, we mark the best and second best results in bold and cyan. For the LLaVA benchmark, we use the LLaVA official code, and for the Qwen2.5-VL-Instruction benchmark. We use lmms-eval (Zhang et al., 2024a) for evaluation.
@@ 5.2 Datasets
In our paper, we mainly divided into 2 scenarios: Image Classification and MLLM Benchmark. The detailed information about datasets is described in Appendix B.1. For the image classification datasets, we choose 5 public classification datasets, including the small-scale dataset of CIFAR100 (Krizhevsky et al., 2009), the large-scale dataset of ImageNet-1K (Russakovsky et al., 2015), and the fine-grained datasets of CUB200 dataset (Wah et al., 2011), FGVC-Aircrafts dataset (Maji et al., 2013), and Stanford-Cars dataset (Krause et al., 2013). For the MLLM datasets, we choose 16 datasets, including visual question answering (VQAv2 (Goyal et al., 2017), GQA (Hudson and Manning, 2019), VizWiz (Gurari et al., 2018), ScienceVQAI (Lu et al., 2022), TextVQA (Singh et al., 2019), MME-RealWorldQA (Zhang et al., 2025c)), understanding (MME (Perception) (Yin et al., 2023), MMBench (Liu et al., 2025), $^{\text{CN}}, MMBench$$^{\text{CC}}$, POPE (F1 score) (Li et al., 2023b), SEEDI (Li et al., 2023a), MMStar (Chen et al., 2024b)), and reasoning (MMMU (Yue et al., 2024a), MMMU-Pro Standard (MMMU-Pros) (Yue et al., 2024b), MathVista (Lu et al., 2024)).